OpenStudy (anonymous):

Help! Did I simplify these equations right? 3x = 5 - 2y, 3x + 2y = 7 simplified to the below.. y = (-3/2)x + (5/2), y = (-3/2)x + 7/2 I'm trying to get these equations into a form, that I can use to put them into the graphing calculator...

OpenStudy (jit4won):

yes

OpenStudy (jit4won):

i have an easier way

OpenStudy (anonymous):

an easier way for what?

OpenStudy (anonymous):

yes if you want to graph them that is what you have to do. because to graph you need y = something

OpenStudy (anonymous):

what you've done is ok indian

OpenStudy (jit4won):

for solving the equation without using the calculator

OpenStudy (anonymous):

I've graphed my simplified equations..but I'm not coming up with the answer that my online test came up with...it says I have the wrong answer...and that there is no solution, it says it's inconsistent. jit4won...my method may work, when I don't have a time limit...but when I have about 7 such problems to do, and I try to work them out on paper, using a normal calculator to figure out the values...and a 1 hour time limit plus other, different problems....I get rushed.

OpenStudy (jit4won):

Do you have skype ?

OpenStudy (anonymous):

Nope, wish I did. Webcams cost money, you know.

OpenStudy (jit4won):

no cam is required

OpenStudy (anonymous):

I don't think there's one built into the family computer here either..

OpenStudy (jit4won):

Do you have anything to hear my voice ?

OpenStudy (anonymous):

indianaelena, did you see what I said on the other post? about the algebraic way to solve systems?

OpenStudy (anonymous):

jitwon, I have no microphone hooked up to the computer either. Good microphones also cost money. Uh, I don't remember seeing that, loophole..

OpenStudy (anonymous):

you might want to go back because the way you are solving takes longer, thats why you couldnt do it on a test...

OpenStudy (anonymous):

I've tried simplifying the equations into y=? form, so I can put them into the calculator, is that what you're talking about, loophole?

OpenStudy (anonymous):

no, go back to your other post, i explained it

OpenStudy (anonymous):

you dont even need a calculator

OpenStudy (jit4won):

okay just read carefully 3x = 5 - 2y means 3x+2y= 5 ................(1) 3x + 2y = 7 ........................................(2) you can see the ratio of coefficients of x = ratio of coefficients of y = 1 but it's not equal to the ratio of constant terms ......... which is 5/7 so it's an inconsistent system and there will be no solution

OpenStudy (anonymous):

So I've looked at your algebraic method, loophole...I've been doing that with past problems...just thought that I could only get the right answer for these by graphing. jit4won, I don't understand your explanation...

OpenStudy (anonymous):

I've tried to simplify the system and graph it on the graphing calculator..and it gave me a solution...

OpenStudy (anonymous):

I must have messed up simplifying it into y=? form..

OpenStudy (anonymous):

So if after simplifying it...the equations are the same...but both come up with different answers, then the system is considered inconsistent?

OpenStudy (jit4won):

yes

OpenStudy (anonymous):

ok...I'm off to get the answers that my online math quiz got...so that hopefully I'll do good on the chapter test, thanks.

OpenStudy (jit4won):

great

OpenStudy (anonymous):

graph of these 2 expressions have same slope but cut the y axis at 5/2 and 7/2 so they are parallel

OpenStudy (jit4won):

yes

OpenStudy (jit4won):

slope are same thats why they are parallel